# Radiative Gaussian Splatting for Efficient X-ray Novel View Synthesis

### v2

@@ 6 Conclusion
In this paper, we propose the first 3DGS-based framework, X-Gaussian, for X-ray novel view synthesis. Our X-Gaussian has two technical components. Firstly, we redesign a radiative Gaussian point cloud model to take the isotropic nature of X-ray imaging into account. This model excludes the influence of view direction when fitting the radiation intensity. Based on this model, we further develop a GPU-friendly differentiable radiative rasterization CUDA kernel that renders projections at a faster speed than RGB rasterization. Secondly, we customize an initialization strategy, ACUI, that does not need to execute the time-consuming SfM algorithm like RGB 3DGS. Instead, ACUI explicitly uses the parameters of X-ray scanner to compute the extrinsic and intrinsic matrices, and then uniformly samples center points for 3D Gaussians within a cuboid enclosing the scanned object. Comprehensive experiments demonstrate that our X-Gaussian significantly outperforms SOTA methods by over 6.5 dB while enjoying shorter training time (15%) and faster inference speed (73$\times$). The application on sparse-view CT reconstruction also shows the potential practical values of our method.

### v3

@@ 6 Conclusion
In this paper, we propose the first 3DGS-based framework, X-Gaussian, for X-ray novel view synthesis. Firstly, we design a radiative Gaussian point cloud model based on X-ray imaging properties. This model excludes the influence of view direction when fitting radiation intensity. For this model, we develop a GPU-friendly differentiable radiative rasterization CUDA kernel that renders projections at a faster speed than RGB rasterization. Secondly, we customize an initialization strategy, ACUI, that does not need to execute the SfM algorithm. Instead, ACUI uses the parameters of X-ray scanner to compute the extrinsic and intrinsic matrices, and then uniformly samples center points for 3D Gaussians within a cuboid enclosing the scanned object. Experiments demonstrate that our X-Gaussian significantly outperforms SOTA methods by over 6.5 dB while enjoying 73$\times$ faster inference speed and only requiring 15% of training time.
6.0.1 Acknowledgements:
This work was supported by the Lustgarten Foundation for Pancreatic Cancer Research and the Patrick J. McGovern Foundation Award.